A force f is expressed with respect to the basis e1 and e2 by equation 2(3e1+4e2) N. If the directions of  e1  makes an angle of 30 degree with f, find the vector  e2.


1
Expert's answer
2021-01-25T13:55:32-0500

Solution

Given force magnitude

f"=6e_1+8e_2"

Angle between e1 and f is =30°

Using basis vector concept

"f. e_1=(6e_1+8e_2). e_1"

"|(6e_1+8e_2) |e_1|cos30^\u00b0=6"

Magnitude of e1 is unit.

So vector e2 can be written as

"e_2=\frac{4\sqrt{3}-6e_1}{8}=\frac{\sqrt{3}}{2}-\frac{3e_1}{4}"